MAG(1) MAG(1) NAME mag−computesfontsizesandmagsteps SYNOPSIS mag[-Rdpi]magstep... mag[-r][-Rdpi]fontsize... DESCRIPTION Thistoolcalculatesfontsizesgivenmagsteps.TeXfontsareprovidedastruesizedfontsoras magnifications.Thefontsizeofatruesizedfontequalstheresolutionoftheprinter(ex.300). Thefontsizeofafontmagnifiednmagstepsequals:1.2n×300roundedtoitsnearestinteger value.FontnamesforTeXfontsnormallyconsistsofthename(cmr),pointsize(10),type(pk) andfontsize(360),forexample:cmr10.360pk. EXAMPLES Theresultofmag-R240-200.51willbe:

167240263288
Theinversefunctioniscomputedwiththe-roption.Theresultofmag-r-R240167
240263288willbethefontsteps:
-200.51
TheUNIXshellsallowcommandsubstitution.Onecanwrite:
mag-r-R240`mag-R240-200.51`

DATE 18December,1989 AUTHOR PietTutelaers UniversityofTechnologyEindhoven rcpt@urc.tue.nl UMFT 1